OpenStudy (anonymous):

find the limit of lim 4x^2-3 over or divide x-7 where x->2/3(two thirds)

OpenStudy (anonymous):

replace x by 2/3 and see what you get. since you do not get 0/0that is your answer.

OpenStudy (anonymous):

if the numerator is 0, and the denominator is not, then there is no limit

OpenStudy (anonymous):

but it isn't so just plug in and see what you get

OpenStudy (anonymous):

? If the numerator is 0 there's no answer? You mean, the denominator is 0? If the numerator is 0, the limit is 0.

OpenStudy (anonymous):

oh i think you should try factoring and i don't know the next step so can you help me..

OpenStudy (anonymous):

yeah of course

OpenStudy (anonymous):

my mistake.

OpenStudy (anonymous):

@moniqu don't factor anything

OpenStudy (anonymous):

replace x by \[\frac{2}{3}\] and see what number you get. that is all

OpenStudy (anonymous):

i get \[\frac{11}{57}\]

OpenStudy (anonymous):

it should be these one:\[(4\chi+3)(4\chi-3) \over \chi-7\]

OpenStudy (anonymous):

ok lets go slow

OpenStudy (anonymous):

you want \[\lim_{x\rightarrow \frac{2}{3}}\frac{4x^2-3}{x-7}\] yes?

OpenStudy (anonymous):

so 11 over 57 is correct?

OpenStudy (anonymous):

this is a rational function and you are looking for the limit at a certain number. now a rational function is continuous everywhere on its domain. the domain of this thing is all numbers except 7, so if you want the limit anywhere else, just plug the number in and see what you get

OpenStudy (anonymous):

i got 11/57 but you should check it. i just computed \[\frac{4(\frac{2}{3})^2-3}{\frac{2}{3}-7}\]

OpenStudy (anonymous):

same... but can I ask on what kind of given should I use factoring.?